# Multi-hop Parallel Image Semantic Communication for Distortion Accumulation Mitigation

### v1

@@ II-B Proposed Multi-hop Parallel Image Semantic Communication Framework
$\mathbf{{\tilde{s}}}_{n}=\mathbf{\hat{s}}_{n}+\mathbf{\hat{r}}_{n}.$ (5)
@@ III-A Network Structure
In accordance to Fig. 1(b), the main network backbone is composed of [7]. The residual compressor is the same as the residual encoder and decoder in [9]. For the residual estimation module, it is shown in Fig. 2. It is composed of a series of residual blocks which mainly contain the convolution layer, generalized divisive normalization layer (GDN), and ReLU layer. After the concatenation through residual blocks, four multi-layer perceptron (MLP) layers finally predict all the required $\mu_{n,i}$, $\sigma_{n,i}$, $\pi_{n,i}$, and $\lambda_{n,i}$ for the discrete mixture of logistic distributions, where $i$ refers to the $i$-th raw image.
@@ 7.     $\mathbf{r}_{n}=\mathbf{s}_{n}-\mathbf{\check{s}}_{n}$
 9.     $\mathbf{{\tilde{s}}}_{n}=\mathbf{\hat{s}}_{n}+\mathbf{\hat{r}}_{n}$
@@ 14.    Update the network weights of $R_{e}(\cdot)$
Figure 3: Quality of the reconstructed images versus the SNRs under Rayleigh fading channels (CBR = 0.071, N = 20).
@@ IV-A4 Evaluation Metrics
We leverage the widely used pixel-wise metric peak signal-to-noise ratio (PSNR), perceptual-level multi-scale structural similarity (MS-SSIM) [15] and learned perceptual image patch similarity (LPIPS) [16] as measurements for the reconstructed image quality.
@@ IV-B1 Performance for Different SNRs
We first evaluate the anti-noise performances of MHPSC under Rayleigh fading channels with a specific CBR, where $CBR=\frac{L+L_{r}}{H\times W\times 3}$. We set the whole CBR as 0.071 while the fixed CBR for the common semantic link is 0.062. As shown in Fig. 3(a), it is clearly to observe that MHPSC outperforms all other benchmarks. Compared to other DL-based schemes, MHPSC outperforms WITT for over 1 dB in terms of PSNR on average, where the gap increases as the SNR value decreases. This trend verifies that parallel link-based MHPSC is more robust to the channel interference than single link-based schemes. With the parallel link, only $13.6\%$ increase of bandwidth cost compared to the original smenatic link enables such performance gain derived from distortion accumulation mitigation. For traditional separated coding schemes, 2/3 code rate LDPC with 16QAM and 1/2 code rate LDPC with 16QAM are set, which reflect corresponding anti-noise performances under different SNR levels. Compared to traditional schemes, MHPSC provides much more performance gain and stability since traditional schemes would be confronted with serious cliff effect with harsh channel conditions. As shown in Fig. 3(b) and Fig. 3(c), the DL-based schemes achieve better reconstruction results in terms of MS-SSIM and LPIPS compared to traditional SSCC schemes, which means that the satisfying visual perception quality is ensured through extracting semantics inside raw images. Compared to other schemes, MHPSC preserves even much more high frequency image details. These results demonstrate the effectiveness of MHPSC for multi-hop transmission under Rayleigh fading channels with various noise intensities.

### v2

@@ II-B Proposed Multi-hop Parallel Image Semantic Communication Framework
$\mathbf{{\tilde{s}}}_{n}=R_{\mathrm{c}_{n}}(\mathbf{\hat{s}}_{n},\mathbf{\hat{r}}_{n}).$ (5)
where $R_{\mathrm{c}_{n}}$ refers to the residual combination network.
@@ III-A Network Structure
In accordance to Fig. 1(b), the main network backbone is composed of [7]. The residual compressor is the same as the residual encoder and decoder in [9]. The residual combination network follows the same U-net structure as [10]. For the residual estimation module, it is shown in Fig. 2. It is composed of a series of residual blocks which mainly contain the convolution layer, generalized divisive normalization layer (GDN), and ReLU layer. After the concatenation through residual blocks, four multi-layer perceptron (MLP) layers finally predict all the required $\mu_{n,i}$, $\sigma_{n,i}$, $\pi_{n,i}$, and $\lambda_{n,i}$ for the discrete mixture of logistic distributions, where $i$ refers to the $i$-th raw image.
@@ 7.     $\mathbf{r}_{n}=\mathbf{s}_{n}-\mathbf{\check{s}}_{n}$
 9.     $\mathbf{{\tilde{s}}}_{n}=R_{\mathrm{c}}(\mathbf{\hat{s}}_{n},\mathbf{\hat{r}}_{n})$
@@ 14.    Update the network weights of $R_{e}(\cdot)$
Figure 3: Quality of the reconstructed images versus the SNRs under Rayleigh fading channels (CBR = 0.077, N = 20).
@@ IV-A4 Evaluation Metrics
We leverage the widely used pixel-wise metric peak signal-to-noise ratio (PSNR), perceptual-level multi-scale structural similarity (MS-SSIM) and learned perceptual image patch similarity (LPIPS) as measurements for the reconstructed image quality.
@@ IV-B1 Performance for Different SNRs
We first evaluate the anti-noise performances of MHPSC under Rayleigh fading channels with a specific CBR, where $CBR=\frac{L+L_{r}}{H\times W\times 3}$. We set the whole CBR as 0.077 while the fixed CBR for the common semantic link is 0.062. As shown in Fig. 3(a), it is clearly to observe that MHPSC outperforms all other benchmarks. Compared to other DL-based schemes, MHPSC outperforms WITT for over 1 dB in terms of PSNR on average, where the gap increases as the SNR value decreases. This trend verifies that parallel link-based MHPSC is more robust to the channel interference than single link-based schemes. With the parallel link, only limited increase of bandwidth cost compared to the original smenatic link enables such performance gain derived from distortion accumulation mitigation. For traditional separated coding schemes, 2/3 code rate LDPC with 16QAM and 1/2 code rate LDPC with 16QAM are set, which reflect corresponding anti-noise performances under different SNR levels. Compared to traditional schemes, MHPSC provides much more performance gain and stability since traditional schemes would be confronted with serious cliff effect with harsh channel conditions. As shown in Fig. 3(b) and Fig. 3(c), the DL-based schemes achieve better reconstruction results in terms of MS-SSIM and LPIPS compared to traditional SSCC schemes, which means that the satisfying visual perception quality is ensured through extracting semantics inside raw images. Compared to other schemes, MHPSC preserves even much more high frequency image details. These results demonstrate the effectiveness of MHPSC for multi-hop transmission under Rayleigh fading channels with various noise intensities.